OpenStudy (anonymous):

how do i do this??? find the % of markup $10.00 marked up to $16.00?

OpenStudy (anonymous):

Well, first I would find the difference between the final amount and the original amount. $16.00 - $10.00 = $6.00 Take this number and divide it by the original amount. $6.00 / $10.00 = 0.6 Multiple by 100 to get the percentage 0.6 * 100 = 60% This is the total percentage markup.

OpenStudy (anonymous):

ok cool so this means 12.50 up to 18.75= 18.75-12.50= 6.25 6.25/12.50=x/100 12.50x=625 12.50x/12.50=625/12.50 X = 50?

OpenStudy (anonymous):

Yes, the best way to check this is to take that 0.5 times the original amount. 0.5 * 12.50 = 6.25 Then add this to the original amount. 12.50 + 6.25 = 18.75 If this number matches the markup number than you did it right!

OpenStudy (anonymous):

sweet thank you!!!!

OpenStudy (anonymous):

Your welcome